Part D -= + HO H2SO4, Draw the molecules on the canvas by choosing buttons from the Tools (for bonds), Atoms, and Advanced TeShow the products.
Answer #1

H₂SO4 +H2O + H₂O 49509 reaction : g - CH3 CHE -GEC-CHE la CH3 - CH₂ - C = C - CH₂ CH₂ - C = C- CH2 H-oso H2o Hot 012 - CH3 -